OpenStudy (anonymous):

Find the midpoint (-6,6) and (-3,-9)

OpenStudy (anonymous):

Midpoint is the middle and middle is the average. (-6,6) and (-3,-9) Average of the x's? (-6 + -3)/2 = -9/2 Average of the y's? (6 + -9)/2 = -3/2 Midpoint is (-9/2, -3/2).

OpenStudy (anonymous):

(x1 + x2)/2, (y1 +y2)/2